a ray of light is refracted through a Prism of angle 60 degree find the angle of incidence so that the emergent ray grazes in the 2nd face.( refractive index of the material of the prism is 1.45°)

Grade:12

1 Answers

Jay vadolkar
5 years ago

Here

=sin{30+Dm/2}/sin60/2

=sin{30+Dm/2}/sin30=√2= 1.414°=>

Sin{30+Dm/2}=1.414×(1/2)=1/√2

=>30+Dm/2=sin-1 (1/√2)=45

=>Dm/2=45–30=15=>Dm= 30°.